# Rolled Gaussian process models for curves on manifolds

### v1

@@ 3.1 Four maps for curves: unrolling, rolling, unwrapping, and wrapping
Conversely, $\gamma$ on $M$ is determined by $\gamma^{\downarrow}$, and we refer thus to $\gamma$ as the rolling of the curve $\gamma^{\downarrow}$ from $T_{\gamma(0)}(M)$ onto $M$. Uniqueness of the parallel transport maps along $\gamma$ in (3.1) (Theorem 4.32, Lee, 2018) ensures that the unrolling and rolling operations are well-defined, and are inverses of each other. Figure 1(a) provides an illustration of the unrolling of a piecewise geodesic curve on $\mathbb{S}^{2}$ onto the tangent space of its initial point. The process of unrolling has an intuitive physical analogy (Jupp and Kent, 1987): imagine that the curve $\gamma:[0,1]\to\mathbb{S}^{2}$ on the unit 2-sphere $\mathbb{S}^{2}$ is drawn in wet ink; its unrolling onto the plane $\mathbb{R}^{2}$, when identified with $T_{\gamma(0)}M$, is the trace of the wet ink left on $\mathbb{R}^{2}$ upon rolling $\mathbb{S}^{2}$ along $\gamma$ on $\mathbb{R}^{2}$ without slipping and twisting. Slipping and twisting is avoided by use of the parallel transport maps, which ensure that the tangent vectors $\dot{\gamma}(t)$ and $\dot{\gamma}^{\downarrow}(t)$ line up at each $t$.
For piecewise geodesic curves, as in the example in Figure 1, a key feature of unrolling, and hence also of the rolling, is that tangent directions and distance between consecutive points along $\gamma$ are preserved in $\gamma^{\downarrow}$, and vice versa. In contrast, a global flattening strategy by projecting the entire curve $\gamma$ onto $T_{\gamma(0)}M$ via the inverse exponential map would distort distances between consecutive points in its image $\exp^{-1}_{\gamma(0)}(\gamma)$ relative to their distances in $\gamma$; there is thus no such distortion associated with the unrolling operation. For modelling purposes, therefore, a mean curve on $M$ may be conveniently defined as the rolling of a curve prescribed on a tangent space.
@@
To define a suitable covariance, we require a sensible notion of deviation of a curve $x$ on $M$ from another fixed curve $\gamma$ on $M$ that is compatible with the rolling/unrolling operations. To begin with, we consider the deviation of a point, $x$, that is associated with with $t\in[0,1]$. Accordingly, given a curve $\gamma:[0,1]\to M$, the unwrapping of point $x$, assumed to be outside the cut locus $\mathcal{C}\{\gamma(t)\}$ of $\gamma(t)$, onto $T_{\gamma(0)}M$ defined as
@@
Let $\mathcal{F}$ and $\mathcal{F}_{b}$ be the space of piecewise smooth curves on $M$ and $T_{b}M$, both with parameter domain $[0,1]$, where $b\in M\backslash\mathcal{C}(\gamma(0))$. Let $\gamma\in\mathcal{F}$, and let $c:[0,1]\to M$ be the geodesic from $b$ to $\gamma(0)$.
@@
    The wrapping map $\mathcal{F}_{b}\to\mathcal{F}$ transforms a curve $y$, with respect to another curve $\gamma\in\mathcal{F}$, into the curve
    $y_{b}^{\uparrow\gamma}(t)=\exp_{\gamma(t)}\left[P^{\gamma}_{t\leftarrow 0}P^{c}_{1\leftarrow 0}\left\{y(t)-\gamma^{\downarrow}_{b}(t)\right\}\right].$ (3.6)
@@
Provided $x$ in (c) is outside the cut loci of $\gamma$, the wrapping map is the unique inverse of the unwrapping map. Provided $y$ in (d) is such that the argument of $\exp_{\gamma(t)}$ in (3.6) is outside the tangential cut locus $\mathcal{T}\{\gamma(t)\}$ for all $t$, then the unwrapping map is the unique inverse of the wrapping map. For manifolds with empty cut loci, unwrapping and wrapping are hence always unique inverses. An example of such a manifold is the set of symmetric positive definite matrices considered in an example in §6.2.
@@
3.2 Rolled Gaussian process model
The maps in the preceding section help define a generative model for random curves on $M$ from a model for curves in $\mathbb{R}^{d}$. Let $z(t)$ be a process in $\mathbb{R}^{d}$ with $E\{z(t)\}=m(t)$, and $y(t)=Uz(t)$ be a process in $T_{b}M$ with $E\{y(t)\}=\tilde{m}(t)=Um(t)$ for an orthonormal frame $U:\mathbb{R}^{d}\rightarrow T_{b}M$. Let the “rolled mean” $\gamma=\tilde{m}^{\uparrow}$ be the rolling of $\tilde{m}$ onto $M$, and $x=z^{\uparrow\gamma}$ be the wrapping of $z$ with respect to $\gamma$. In some circumstances, the rolled mean and Fréchet mean of $x$ coincide.
@@ 3.2 Rolled Gaussian process model
For every $t$, the rolled mean, $\gamma(t)$, is the Fréchet mean of $x(t)$ if any of the following conditions are true:
@@
    Every point in $M$ has an empty cut locus;
@@
    $M$ is a symmetric manifold, the distribution of $x(t)$ has even symmetry about $\gamma(t)$, and has a unique Fréchet mean that lies outside the cut locus of $\gamma(t)$.
Theorem 3.3 holds for general processes $z$ with finite mean, not necessarily Gaussian, and is used later when establishing consistency of the estimator of the $\tilde{m}$ in Theorem 4.2. Condition $(i)$ in Theorem 3.3 applies, for example, to non-positively curved manifolds, e.g. $d$-dimensional hyperbolic space $\mathbb{H}^{d}$ with the hyperbolic metric and the space $\text{Sym}_{>0}(d)$ of $d\times d$ real symmetric positive definite matrices with several choices of Riemannian metrics; examples of symmetric manifolds relevant to condition $(ii)$ are $\mathbb{H}^{d}$ and $\text{Sym}_{>0}(d)$, compact manifolds such as $\mathbb{S}^{2}$, rotation matrices, and Lie groups with bi-invariant metrics.
@@
Given $b\in M$ and a curve $\gamma$ on $M$, let $U=\{u_{1},\ldots,u_{d}\}$ be an orthonormal basis on $T_{b}M$ in its matrix representation. With $\tilde{m}(t)=Um(t):=\sum_{i=1}^{d}m^{i}(t)u_{i}$ and $\tilde{K}(t,t^{\prime})=UK(t,t^{\prime})U^{\top}$, consider the process $y_{b}(t):=Uz(t)$. On the inner product space $(T_{b}M,\langle\cdot,\cdot\rangle_{b})$ note that $y_{b}(t)$ is a Gaussian random vector with mean $\tilde{m}(t)$ and covariance determined by the nonnegative linear operator $\tilde{K}(t,t):T_{b}M\to T_{b}M$; see Appendix 10.1 for details on Gaussians on inner product spaces. As a consequence, $y_{b}(t):=Uz(t)\sim\mathcal{GP}(\tilde{m},\tilde{K})$ assuming values in $T_{b}M$. We can thus identify a Gaussian process in $T_{b}M$ with a Gaussian process $z\sim\mathcal{GP}(m,K)$ in $\mathbb{R}^{d}$, that when transformed using the wrapping map in Definition 3.2(b) with respect to a curve $\gamma$ in $M$, results in a random curve on $M$ whose law is determined by the pair $(b,U)$ and parameters $(m,K)$.
@@
Let $z\sim GP(m,k)$, and choose $b\in M$ and basis $U$ of $T_{b}M$. With $\tilde{m}_{b}=Um$ as a curve in $T_{b}M$, let $\gamma:=\tilde{m}_{b}^{\uparrow}$ be the rolling of $\tilde{m}$, and let $x=y_{b}^{\uparrow\gamma}$ be the wrapping of $y_{b}:=Uz$ with respect to $\gamma$. The random curve $x$ on $M$ is a rolled Gaussian process, denoted $x\sim\mathcal{RGP}(m,k;b,U)$.
Of particular note in the definition above is the fact that the Gaussian process $z$ in $T_{b}M$ is wrapped with respect to the rolling of its mean function $\tilde{m}$. This is a convenient modelling feature because the curve $\gamma$ then is not an additional parameter to be estimated, but is instead fully determined by the mean of $z$. Figure 2 illustrates the construction of the rolled Gaussian process and notation.
@@
The result below identifies the rolled Gaussian process $x$ as the image of a Gaussian vector field along $\gamma$ under the exponential map, and may be compared with the construction in Mallasto and Feragen (2018). It uses the notion of a frame along $\gamma$ detailed in the Appendix 10.2.
@@
Let $E^{\gamma}=\{e_{1},\ldots,e_{d}\}$ be a frame along $\gamma$. For the pair $(b,U)$, the random vector field $t\mapsto V^{\gamma}(t):=E^{\gamma}(t)y_{b}(t)=\sum_{i=1}^{d}y_{b}^{i}(t)e_{i}(t)\in T_{\gamma(t)}M$ along $\gamma$ is a Gaussian vector field.
The choice of point $b\in M$ and basis $U$ on its tangent space $T_{b}M$ is arbitrary. The rolled Gaussian process, however, is equivariant with respect to this choice.
@@
A consequence of this equivariance is that from a practical modelling and computational point of view, the particular choice of $b$ and $U$ is inconsequential.
3.3 A parametric Gaussian process model for random curves on $\mathbb{R}^{d}$
A core ingredient of the rolled Gaussian process model in the preceding section is a Gaussian process model, $z\sim\mathcal{GP}(m,K)$, for curves in $\mathbb{R}^{d}$ to be rolled and wrapped onto $M$. This model requires specification of the mean, $m$, and covariance, $K$, functions. Furthermore, for practical calculations it is helpful to introduce a basis representation for curves, in terms of which $m$ and $K$ can be prescribed and estimated parametrically.
Let $\{\phi_{s}:[0,1]\rightarrow\mathbb{R}\}$ be a B-spline basis (De Boor, 2002; Ramsay et al., 2009) of order 4 and knots at the $k-2$ interior points of a grid of $k$ equally spaced points on $[0,1]$ with $k\geq\operatorname{max}(4,d)$; then each $\phi_{s}$ is a piecewise cubic polynomial with joins at the knots such that any adjacent polynomials’ values, and values of first and second derivatives, are matched (De Boor, 2002). By construction, $\sum_{s=1}^{k}\phi_{s}(t)=1$, for any $t\in[0,1]$; that is, the constant function is in the span of this basis, a property useful later in Proposition 4.1. Let the mean $m(t)=M_{w}\phi(t)$, and assume a separable covariance $K(t,t^{\prime})=\phi(t)^{\top}V_{w}\phi(t^{\prime})U_{w}$, where $\phi(t)=\{\phi_{1}(t),\ldots,\phi_{k}(t)\}\in\mathbb{R}^{k}$ is a vector and $M_{w}\in\mathbb{R}^{d\times k}$, $U_{w}\in\operatorname{Sym}_{>0}(d)$ and $V_{w}\in\operatorname{Sym}_{>0}(k)$ are matrices that parameterise the model, where $\operatorname{Sym}_{>0}(p)$ is the cone of $p\times p$ symmetric positive definite matrices. The convenience of this particular choice is that the curve $z$ can be written

### v2

@@ 3 The unrolling, rolling, unwrapping, and wrapping maps
Conversely, $\gamma$ on $M$ is determined by $\gamma^{\downarrow}$, and we refer thus to $\gamma$ as the rolling of the curve $\gamma^{\downarrow}$ from $T_{\gamma(0)}(M)$ onto $M$. Uniqueness of the parallel transport maps along $\gamma$ in (3.1) (Theorem 4.32, Lee, 2018) ensures that the unrolling and rolling operations are well-defined, and are inverses of each other. Figure 1(a) provides an illustration of the unrolling of a piecewise geodesic curve on $\mathbb{S}^{2}$ onto the tangent space of its initial point. The process of unrolling has an intuitive physical analogy (Jupp and Kent, 1987): imagine that the curve $\gamma:[0,1]\to\mathbb{S}^{2}$ on the unit 2-sphere $\mathbb{S}^{2}$ is drawn in wet ink; its unrolling onto the plane $\mathbb{R}^{2}$, when identified with $T_{\gamma(0)}M$, is the trace of the wet ink left on $\mathbb{R}^{2}$ upon rolling $\mathbb{S}^{2}$ along $\gamma$ on $\mathbb{R}^{2}$ without slipping. Slipping is avoided by use of the parallel transport maps, which ensure that the tangent vectors $\dot{\gamma}(t)$ and $\dot{\gamma}^{\downarrow}(t)$ line up at each $t$. When $\mathbb{S}^{2}$ is viewed as an embedded submanifold of $\mathbb{R}^{3}$, the extra dimension demands that in addition to no slipping one also avoids “twisting” (Sharpe, 2000). Thus, it is appropriate to view $\gamma$ and $\gamma^{\downarrow}$ as two equivalent copies of the same curve.
For piecewise geodesic curves, as in the example in Figure 1, a key feature of unrolling, and hence also of the rolling, is that segment lengths and angles between consecutive segments along $\gamma$ are preserved in $\gamma^{\downarrow}$, and vice versa. In contrast, a global flattening strategy by projecting the entire curve $\gamma$ onto $T_{\gamma(0)}M$ via the inverse exponential map would distort these angles and lengths in its image $\exp^{-1}_{\gamma(0)}(\gamma)$ relative those in $\gamma$. There is thus no such distortion when unrolling. For modelling purposes, therefore, a mean curve on $M$ may be conveniently defined as the rolling of a curve prescribed on a tangent space.
@@
To define a suitable covariance, we require a notion, compatible with rolling/unrolling, of how a curve $x$ on $M$ deviates relative to another fixed curve $\gamma$ on $M$. To begin with, we consider the deviation of a point, $x$, that is associated with with a particular $t\in[0,1]$. Accordingly, relative to $\gamma(t)$, the unwrapping of point $x$, assumed to be outside the cut locus $\mathcal{C}\{\gamma(t)\}$ of $\gamma(t)$, onto $T_{\gamma(0)}M$ defined as
@@
Let $\mathcal{F}$ and $\mathcal{F}_{b}$ be the space of piecewise smooth curves on $M$ and $T_{b}M$, both with parameter domain $[0,1]$, where $b\in M\backslash\mathcal{C}(\gamma(0))$. Let $\gamma\in\mathcal{F}$, and let $c:[0,1]\to M$ be a geodesic from $b$ to $\gamma(0)$.
@@
    The wrapping map $\mathcal{F}_{b}\to\mathcal{F}$ transforms a curve $y_{b}$, with respect to another curve $\gamma\in\mathcal{F}$, into the curve
    $y_{b}^{\uparrow\gamma}(t)=\exp_{\gamma(t)}\left[P^{\gamma}_{t\leftarrow 0}P^{c}_{1\leftarrow 0}\left\{y_{b}(t)-\gamma^{\downarrow}_{b}(t)\right\}\right].$ (3.6)
@@
Provided $x$ in (c) is outside the cut loci of $\gamma$, the wrapping map is the unique inverse of the unwrapping map. Provided $y_{b}$ in (d) is such that the argument of $\exp_{\gamma(t)}$ in (3.6) is outside the tangential cut locus $\mathcal{T}\{\gamma(t)\}$ for all $t$, then the unwrapping map is the unique inverse of the wrapping map. For manifolds with empty cut loci, unwrapping and wrapping are hence always unique inverses. An example of such a manifold is the set of symmetric positive definite matrices considered in an example in §8.2.
@@
4 Rolled Gaussian process model
The maps in the preceding section help define a generative model for random curves on $M$ from a model for curves in $\mathbb{R}^{d}$. Let $z(t)$ be a process in $\mathbb{R}^{d}$ with $E\{z(t)\}=m(t)$, and $y_{b}(t)=Uz(t)$ be a process in $T_{b}M$ with $E\{y_{b}(t)\}=m_{b}(t)=Um(t)$ for an orthonormal frame $U:\mathbb{R}^{d}\rightarrow T_{b}M$. Let the “rolled mean” $\gamma=m_{b}^{\uparrow}$ be the rolling of $m_{b}$ onto $M$, and $x=y_{b}^{\uparrow\gamma}$ be the wrapping of $y_{b}$ with respect to $\gamma$. In terms of notation, for a fixed $b\in M$:
  • $\bullet$
    A subscript $b$ without a superscript distinguishes fixed or random curves and parameters on $T_{b}M$;
  • $\bullet$
    $z$ denotes an $\mathbb{R}^{d}$-valued stochastic process, and $x$ an $M$-valued process.
The following result elucidates on conditions under which the rolled mean $m_{b}^{\uparrow}$ equals the population Fréchet mean $\gamma_{\text{Fre}}$ of $x$.
@@ 4 Rolled Gaussian process model
Assume that for every $t\in[0,1]$, the Fréchet mean of $x(t)$ exists and is unique. Then, for every $t$, the rolled mean, $\gamma(t)$, is the Fréchet mean of $x(t)$ if any of the following conditions are true:
@@
    $M$ is a Hadamard manifold;
@@
    $M$ is a symmetric space, the distribution of $x(t)$ has geodesic symmetry about $\gamma(t)$, and has a unique Fréchet mean that lies outside the cut locus of $\gamma(t)$.
Theorem 4.1 holds for general processes $z$ in $\mathbb{R}^{d}$ with finite mean, not necessarily Gaussian. It plays an important role in §5 where estimators of the mean function $m$ and covariance $k$ of a rolled Gaussian process are proposed and their asymptotic properties are studied. Manifolds in (i) of Theorem 4.1 include the $d$-dimensional hyperbolic space $\mathbb{H}^{d}$ with the hyperbolic metric and the space $\text{Sym}_{>0}(d)$ of $d\times d$ real symmetric positive definite matrices with several choices of Riemannian metrics; examples of symmetric manifolds relevant to condition (ii) are the non-compact $\mathbb{H}^{d}$ and $\text{Sym}_{>0}(d)$, and compact manifolds such as $\mathbb{S}^{2}$, rotation matrices, and Lie groups with bi-invariant metrics.
@@
Given $b\in M$ and a curve $\gamma$ on $M$, let $U=\{u_{1},\ldots,u_{d}\}$ be an orthonormal basis on $T_{b}M$ in its matrix representation. With $m_{b}(t)=Um(t):=\sum_{i=1}^{d}m^{i}(t)u_{i}$ and $K_{b}(t,t^{\prime})=UK(t,t^{\prime})U^{\top}$, consider the process $y_{b}(t):=Uz(t)$. On the inner product space $(T_{b}M,\langle\cdot,\cdot\rangle_{b})$ note that $y_{b}(t)$ is a Gaussian random vector with mean $m_{b}(t)$ and covariance determined by the nonnegative linear operator $K_{b}(t,t):T_{b}M\to T_{b}M$; see Appendix A.1 for details on Gaussians on inner product spaces. As a consequence, $y_{b}(t):=Uz(t)\sim\mathcal{GP}(m_{b},K_{b})$ assuming values in $T_{b}M$. We can thus identify a Gaussian process in $T_{b}M$ with a Gaussian process $z\sim\mathcal{GP}(m,K)$ in $\mathbb{R}^{d}$, that when transformed using the wrapping map in Definition 3.2(b) with respect to a curve $\gamma$ in $M$, results in a random curve on $M$ whose law is determined by the pair $(b,U)$ and parameters $(m,K)$.
@@
Let $z\sim GP(m,K)$, and choose $b\in M$ and a frame $U:\mathbb{R}^{d}\to T_{b}M$. With $m_{b}=Um$ as a curve in $T_{b}M$, let $\gamma:=m_{b}^{\uparrow}$ be the rolling of $m_{b}$, and let $x=y_{b}^{\uparrow\gamma}$ be the wrapping of $y_{b}:=Uz$ with respect to $\gamma$. The random curve $x$ on $M$ is a rolled Gaussian process, denoted $x\sim\mathcal{RGP}(m,K;b,U)$.
Of particular note in the definition above is the fact that the Gaussian process $y_{b}$ in $T_{b}M$ is wrapped with respect to the rolling of its mean function $m_{b}$. This is a convenient modelling feature because the curve $\gamma$ then is not an additional parameter to be estimated, but is instead fully determined by the mean of $z$. Figure 2 illustrates the construction of the rolled Gaussian process and notation.
@@
The rolled Gaussian process involves a random vector field along $\gamma$ arising from the wrapping operation. Proposition 4.1 below identifies it as a Gaussian vector field, which we claim is a natural one. To understand why, we note that if Gaussian random variables $v(p)\sim N_{d}(\mu(p),\Sigma(p))$ on $T_{p}M$ and $v(q)\sim N_{d}(\mu(q),\Sigma(q))$ on $T_{q}M$ are such that
$\mu(q)=P^{\gamma}_{0\to 1}\mu(p),\quad\Sigma(q)=P^{\gamma}_{0\to 1}\Sigma(p)P^{\gamma}_{0\leftarrow 1},$
where $\gamma:[0,1]\to M$ is a piecewise smooth curve taking $\gamma(0)=p$ to $\gamma(1)=q$. Then, $v(q)\overset{d}{=}P^{\gamma}_{0\to 1}v(p)$ and it is easy to see using geodesic normal coordinates that $v(q)$ and $v(p)$ are Gaussian random vectors on two isometric copies of $\mathbb{R}^{d}$. Back to our setting, consider the random field $t\mapsto V^{\gamma}(t)\in T_{\gamma(t)}$, $V^{\gamma}(t):=P^{\gamma}_{t\leftarrow 0}P^{c}_{1\leftarrow 0}(y_{b}(t)-m_{b}(t))$. The tangent vectors $V^{\gamma}(s)\in T_{\gamma(s)}M$ and $V^{\gamma}(t)\in T_{\gamma(t)}M$ are related via parallel transport maps, and each a zero-mean Gaussian since parallel transports are linear isometries. Thus $V^{\gamma}$ along $\gamma$ may be viewed as an isometric copy of the $\mathbb{R}^{d}$-valued Gaussian process $z$. Adapting terminology from the deterministic setting, we say that the vector field $V^{\gamma}$ is covariantly Gaussian, and in this sense, the rolled Gaussian process is the natural curved analogue of the Euclidean Gaussian process.
@@
Given $(b,U)$, $\{V^{\gamma}(t),t\in[0,1]\}$ is a covariant Gaussian vector field along $\gamma$.
The choice of point $b\in M$ and basis $U$ on its tangent space $T_{b}M$ is arbitrary. The rolled Gaussian process, however, is equivariant with respect to this choice so that in practice the particular choice of $b$ and $U$ is inconsequential.
@@
4.1 Parametric models
A core ingredient of the rolled Gaussian process model is a Gaussian process model, $z\sim\mathcal{GP}(m,K)$, for curves in $\mathbb{R}^{d}$ to be rolled and wrapped onto $M$. This model requires specification of the mean, $m$, and covariance, $K$, functions. For many applications it is beneficial to consider $z$ in a basis representation, which may be defined as follows. Let $\{\phi_{s}:[0,1]\rightarrow\mathbb{R}\}$ be a linearly independent basis for real functions on $[0,1]$, which includes the constant function. Let $z(t)=\sum_{s=1}^{\infty}\xi_{s}\phi_{s}(t)$, where $\xi=(\xi_{1},\xi_{2},\ldots)$ is a random Gaussian element in a Hilbert space $H$. Its mean $\mathbb{E}z(t)=m(t)=K_{0}\Phi(t)$, where $\Phi(t)=(\phi_{1}(t),\phi_{2}(t),\ldots)\in\ell^{2}$, and $K_{0}:\ell^{2}\to\mathbb{R}^{d}$ is a linear map; its covariance $\Xi:H\times H$ given by the tensor product $\Xi=V\otimes U$, where $U\in\text{Sym}_{>0}(d)$ and $V:\ell^{2}\to\ell^{2}$ is a bounded, self-adjoint operator on the space $\ell^{2}$ of square-summable real sequences. Thus the associated covariance function $K(s,t)=\langle\Phi(s),V\Phi(t)\rangle_{H}$U.
The special case with a finite number, $k$, of basis functions is particularly useful when data curves are observed at discrete times, as in the robot application in §9. For example, let $\{\phi_{s}:[0,1]\rightarrow\mathbb{R}\}$ be a B-spline basis (De Boor, 2002; Ramsay et al., 2009) of order 4 and knots at the k2k-2 interior points of a grid of $k$ equally spaced points on $[0,1]$ with $k\geq\operatorname{max}(4,d)$; then each $\phi_{s}$ is a piecewise cubic polynomial with joins at the knots such that any adjacent polynomials’ values, and values of first and second derivatives, are matched (De Boor, 2002). By construction, $\sum_{s=1}^{k}\phi_{s}(t)=1$, for any $t\in[0,1]$; that is, the constant function is in the span of this basis, a property useful later in Proposition 6.1. The Gaussian process $z$ can be written as